# UCIL: An Unsupervised Class Incremental Learning Approach for Sound Event Detection

### v1

@@ 3.2 Implementation Details and Reference Baselines
4 Results and Analysis
4.1 Results for Two-task and Four-Task Settings
We first conduct the studies under the two-task setting on the DCASE 2023 Task 4A dataset, whose results are reported in Table 1. We consider random sampling to add the rehearsal data in EWC and LwF methods for a fair comparison. We focus on demonstrating how CIL minimizes the performance gap due to catastrophic forgetting. It observed that the lower-bound finetune approach faces challenges in adapting to new tasks without retention strategies as expected. The PSDS score is observed to be zero which indicates the model forgot most of the previous knowledge and similar observation was also reported in [20]. Analyzing the impact of different CIL strategies between scenarios, it is evident that each method’s ability to handle the nuances of prompt detection (PSDS1) versus class confusion (PSDS2) varies. Our UCIL method, particularly with the highest rehearsal size, shows notable performance, closely following the upper bound in both scenarios. This highlights the competence of UCIL in rapidly reacting to new sounds while effectively managing class distinctions. The performance of UCIL method across different rehearsal sizes also provides insights into its scalability. While a decrease in rehearsal size leads to a slight reduction in PSDS, UCIL maintains competitive performance. This is particularly significant on comparing UCIL to other CIL approaches such as EWC, LwF, and NR, where UCIL consistently outperforms or matches their scores, specifically for PSDS2.
Table 2: Performance comparison of our proposed UCIL method with other CIL methods using the four-task settings. Rehearsal size is varied similar to two-task setting.
@@ 4.1 Results for Two-task and Four-Task Settings
We then conduct the studies for four-task setting that are reported in Table 2. It is observed that UCIL demonstrates a remarkable performance, especially notable in the PSDS2 scores, where its ability to minimize class confusion gets highlighted. With a rehearsal size of [2000, 200], UCIL’s PSDS2 score doubles that of NR significantly outperforms other CIL methods like EWC and LwF. Moreover, the comparison between UCIL’s performance with different rehearsal sizes provides valuable insights into its efficiency. Despite a reduction in the amount of strongly labeled data, UCIL maintains commendable performance, showcasing its effective use of data under a higher task number based CIL studies.
Table 3: Ablation studies for UCIL method with different settings. All experiments are with [1000, 200] rehearsal data size.
Method FD UL MU PSDS1 PSDS2
UCIL 0.184 0.332
UCIL 0.081 0.112
UCIL 0.073 0.143
UCIL 0.110 0.193
4.2 Ablation Study
Table 3 shows an ablation study of our UCIL method, examining the significance of each components under the four-task setting. Each experiment is conducted with a fixed rehearsal size of [1000, 200] for strongly and weakly labeled data, providing a controlled environment to isolate the impact of each module: feature distillation (FD), unsupervised learning (UL), and memory update (MU). We find that the best performance is achieved when all components are active (FD, UL, MU). Removing MU leads to a notable drop in scores, highlighting MU’s critical role in effectively integrating new information while retaining previous knowledge. Similarly, excluding UL also decreases performance, underscoring UL’s importance in leveraging unlabeled data for broader learning. Excluding FD while keeping UL and MU active shows a lesser but still notable decline in performance, highlighting FD’s contribution to maintaining model consistency and effectiveness across varied sound events.
5 Conclusion
In this work, we proposed a novel method for CIL referred to as UCIL to advance SED research. Incorporating strategies such as independent learning, temporal knowledge distillation and unsupervised sample selection, our method stands out for its ability to maintain model consistency and leverage unlabeled data effectively. The balanced update mechanism further ensures equitable representation of sound classes. The studies on the DCASE 2023 Task 4 dataset, not only showcases the practical applicability of UCIL, but also sets a stage for future research in dynamic audio environments.

### v2

@@ III-B Implementation Details and Reference Baselines
IV Results and Analysis
IV-A Results for Two-task and Four-Task Settings
We first conduct the studies under the two-task setting on the DCASE 2023 Task 4A dataset, whose results are reported in Table I. Rehearsal size indicates the maximum clip numbers of [Strongly Labeled, Weakly Labeled] as the rehearsal data in each task. We consider random sampling to add the rehearsal data in EWC and LwF methods for a fair comparison. We focus on demonstrating how CIL minimizes the performance gap due to catastrophic forgetting. It observed that the lower-bound finetune approach faces challenges in adapting to new tasks without retention strategies as expected. The PSDS score is observed to be zero which indicates the model forgot most of the previous knowledge and similar observation was also reported in [22]. Analyzing the impact of different CIL strategies between scenarios, it is evident that each method’s ability to handle the nuances of prompt detection (PSDS1) versus class confusion (PSDS2) varies. Our UCIL method, particularly with the highest rehearsal size, shows notable performance, closer to the upper bound in both metrics. This highlights the competence of UCIL in rapidly reacting to new sounds while effectively managing class distinctions. The performance of UCIL method across different rehearsal sizes also provides insights into its scalability. While a decrease in rehearsal size leads to a slight reduction in PSDS, UCIL maintains competitive performance. This is particularly significant on comparing UCIL to other CIL approaches such as EWC, LwF, and NR, where UCIL consistently outperforms or matches their scores, specifically for PSDS2.
TABLE II: Performance comparison of our proposed UCIL method with other CIL methods using the four-task settings. Rehearsal size is varied similar to two-task setting.
@@ IV-A Results for Two-task and Four-Task Settings
We then conduct the studies for four-task setting that are reported in Table II. It is observed that UCIL demonstrates a remarkable performance, especially notable in the PSDS2 scores, where its ability to minimize class confusion gets highlighted. With a rehearsal size of [2000, 200], UCIL’s PSDS2 score doubles that of NR and significantly outperforms other CIL methods like EWC and LwF. Moreover, the comparison between UCIL’s performance with different rehearsal sizes provides valuable insights into its efficiency. Despite a reduction in the amount of strongly labeled data, UCIL maintains commendable performance, showcasing its effective use of data under a higher task number based CIL studies.
TABLE III: Ablation studies for UCIL method with different settings. All experiments are with [1000, 200] rehearsal data size.
FD UL MU PSDS1 PSDS2
0.184 0.332
0.081 0.112
0.073 0.143
0.110 0.193
w/o Selection 0.171 0.284
w/o Balance 0.135 0.217
IV-B Ablation Study
Table III shows an ablation study of our UCIL method, examining the significance of each component under the four-task setting. Each experiment is conducted with a fixed rehearsal size of [1000, 200] for strongly and weakly labeled data, providing a controlled environment to isolate the impact of each module: feature distillation (FD), unsupervised learning (UL), and memory update (MU). We find that the best performance is achieved when all components are active (FD, UL, MU). Removing MU or even leads to a notable drop in scores, highlighting MU’s critical role in effectively integrating new information while retaining previous knowledge. Similarly, excluding UL also decreases performance, underscoring UL’s importance in leveraging unlabeled data for broader learning. Excluding FD while keeping UL and MU active shows a lesser but still notable decline in performance, highlighting FD’s contribution to maintaining model consistency and effectiveness across varied sound events. Additionally, we observe that removing the (w/o) selection process for unlabeled data results in a moderate drop in performance, indicating the importance of selection. The model struggles more when removing (w/o) balance sampling, suggesting that balanced rehearsal data maintains performance across all sound classes.
V Conclusion
In this work, we proposed a novel method for CIL referred to as UCIL to advance SED research. Incorporating strategies such as independent learning, knowledge distillation and unsupervised sample selection, our method stands out for its ability to maintain model consistency and leverage unlabeled data effectively. The balanced update mechanism further ensures equitable representation of sound classes. The studies on the DCASE 2023 Task 4 dataset, not only showcase the practical applicability of UCIL, but also set a stage for future research in dynamic audio environments.